# expected utility

### v2

@@ 2.2 Optimization Functional
where $X_{s}=x$, $0\leq k_{0},k_{1}\leq k_{2}<\infty$ with $k_{0}+k_{1}\leq k_{2}$, $0\leq\alpha_{2}<\alpha<\infty$ with $\tilde{\alpha}:=\alpha-\alpha_{2}$. Note that $0<\tilde{\alpha}\leq\alpha$ holds. We also write $F(X_{T})$ instead of $F(0,T,u,x)$ when the trading strategy is clear (for instance, for the optimal terminal wealth $\hat{X}_{T}$), suppressing the initial value.
@@
Since the problem in this paper is a variant of a mean-variance optimization, it inherits also the time-inconsistency of the classical mean-variance approach, see, for instance, Basak and Chabakauri [3] or Björk and Murgoci [5]. There are several ways to approach this time-inconsistency. We take in this paper the perspective of a so-called pre-committed investor. Such an investor optimizes once in the beginning and sticks to this strategy, i.e., the chosen strategy of the investor is not dynamically reoptimized. Note that due to the function $F$, our problem is not covered by the general theory from Björk and Murgoci [5] which assumes a sophisticated investor, i.e., she/he knows that her/his optimal strategy changes over time and only optimizes over all strategies she/he would actually follow.
@@ 3.1 Derivation of the optimal terminal wealth
    The representation of $\hat{X}_{T}$ also holds in a general complete market (without necessarily using a Black-Scholes market model) if we assume the existence of the multipliers $y$ and $\lambda$. However, that these exist is then not clear. Note that the price density process, $\xi$, in a different complete market would change, but still admit a geometric form. For example, if we replace the Brownian Motion $W$ by a Poisson process $N$ with parameter $\lambda_{2}$ (in the Black-Scholes model), the price density admits a dynamic of the form (see, e.g., Cont and Tankov [12, p.303]):
@@
    The following paragraph considers $L$ as a function of $X$. Obviously, $L$ is not smooth in $k_{1}$ and $k_{2}$, but it is continuous in $[0,\infty)$. Hence, we optimize $L$ in the regions $[0,k_{1}]$, $[k_{1},k_{2}]$, and $[k_{2},\infty)$ separately and compare afterwards the minimal points. Now, for the individual optimization, we get:
@@
    $X_{2}\in[k_{1},k_{2}]$$\Leftrightarrow$ $\xi\in\left[\displaystyle\frac{\lambda\alpha-2\gamma\alpha^{2}(k_{2}-k_{1}-k_{0})}{y},\displaystyle\frac{\lambda\alpha}{y}+\displaystyle\frac{2\gamma\alpha^{2}k_{0}}{y}\right]=\left[\alpha\hat{\xi},\bar{\xi}\right],$ (3.6)
@@
    We notice that $\xi_{+}>\bar{\xi}$ and $\max\left\{\alpha\hat{\xi},\xi_{-}\right\}=\xi_{3}^{*}$ where $\xi_{3}^{*}$ is defined as in Theorem 3.2. If $\xi_{-}<\alpha\hat{\xi}$, $X_{2}$ is not in the necessary interval $[k_{1},k_{2})$ which entails that $(\alpha\hat{\xi},\xi_{3}^{*}]=\emptyset$. Otherwise, it holds that $\xi_{-}=\xi_{3}^{*}$ by which we get (3.4) for this case, i.e., the interval for $\xi$ being in the interval $(\alpha\hat{\xi},\xi_{3}^{*}]$ for the optimal wealth $\hat{X}_{T}$.
@@
    Then it holds:
@@
    Again, we notice that $\xi_{+}>\tilde{\alpha}\hat{\xi}$ and if the discriminant is non-negative, we get $\xi_{-}=\tilde{\xi}_{1}^{*}$ where we can ignore this case if $\xi_{-}\leq 0$ due to the assumption that $\xi$ has to be in the interval $(0,\tilde{\alpha}\hat{\xi}]$. Now, if the discriminant is negative, then $L(X_{3},y,\xi)\geq-\gamma\alpha^{2}k_{0}^{2}-\lambda\alpha k_{0}$ for all $\xi$ and $\tilde{\xi}_{1}^{*}>\tilde{\alpha}\hat{\xi}$. Hence, $X_{3}$ is optimal for all $\xi\in(0,\tilde{\alpha}\hat{\xi}]$ and it holds that $\xi_{1}^{*}=\tilde{\alpha}\hat{\xi}$ which implies (3.4) for this case.
@@
    Hence, it follows that $\xi\in(\tilde{\alpha}\hat{\xi},\tilde{\xi}_{2}^{*}]$ which implies (3.4) in this case as well. Hence, (3.4) holds in total.
@@ 4 Numerical Results
In Figure 4, we present the optimal terminal wealth as a function of $\xi_{T}$ for different values of risk aversion parameter $\gamma$. (Note that the risk aversion parameters are not comparable for the different optimization strategies.) For both the non-protected and the protected case, we compare the optimal terminal wealth stemming from our mean-variance optimization with the EU-optimization for participating life insurance contracts from Lin et al. [25]. To measure the utility, we use, as Lin et al., the following S-shaped utility function: $U(x)=\begin{cases}x^{\tilde{\gamma}}&x\geq 0,\\ -\tilde{\lambda}(-x)^{-\tilde{\gamma}}&x<0\end{cases}$ with $\tilde{\lambda}=2$ and different values for $\tilde{\gamma}$ (Lin et al. [25] used $\tilde{\gamma}=0.5$ in their numerical analysis). Moreover, we added the case that there is no additional surplus participation of the policyholders, i.e., $\alpha_{2}=0$, $k_{0}=0$, and $k_{1}=0$. Both participating optimizations share that the optimal wealth function has three points of significant behavioral change, i.e., at $\xi_{1}^{*}$, $\xi_{2}^{*}$ and $\xi_{3}^{*}$ as in Theorem 3.2. This theorem shows that the optimal wealth decreases before $\xi_{1}^{*}$ and between $\xi_{2}^{*}$ and $\xi_{3}^{*}$, and that the optimal wealth is constant elsewhere. The main difference in the shape of the functions is that we have a piecewise linearity in the mean-variance case but a total non-linearity in the EU case. This effect is already apparent from the respective formulas. Noticeably, the distance between the second and third point of behavioral change is much higher in the EU-optimization than in the mean-variance one. Moreover, the figure shows a moral hazard of the insurance companies offering a non-protected participating life insurance contract since the insurance companies favor a portfolio value of 00 over a portfolio value at the guarantee value level. We can observe this effect for both the mean-variance and the EU-optimization in the figure due to the drop-down of the optimal terminal wealth to zero. Protected products have no such effect when optimizing mean-variance since the insurance company always profits from higher portfolio values. While this effect also occurs when optimizing EU, the probability is smaller since the drop-down is at a far higher value (compared to the non-protected product).
@@
In both, Figure 4 and Figure 5, we can observe a jump of the optimal terminal wealth in the non-protected case, which is typical for problems with option-like constraints, see for instance Basak and Shapiro [4], Chen et al. [8], Nguyen and Stadje [33], Chen et al. [9], or Kraft and Steffensen [21]. The reason is that for certain outcomes, the insurance company’s incentive change and therefore it can gain additional utility by changing to a different regime. [^6] For instance, if the entire wealth for certain outcomes is transferred to the policyholders, the insurance company will target a terminal wealth of zero in these states in order to gain additional wealth in states where it benefits more. (Note that the overall expectation is fixed due to the budget constraint.) Moreover, the piecewise linearity observed also occurs in, e.g., Avanzi et al. [1] who considered a minimization of a quadratic function under solvency constraints including option-like constraints as the expected shortfall. This piecewise linearity is typical for problems with some kind of quadratic penalties, whereas the option-like constraints entail the existence of the plateaus and the jumps.
@@ Appendix A Proofs
    We start with the proof of (i). Therefore, we reformulate the condition $\xi_{3}^{*}>\alpha\hat{\xi}$ to get a condition for $\lambda$. When plugging in the formulas as defined in Theorem 3.2, it holds that $\xi_{3}^{*}>\alpha\hat{\xi}$ if and only if
@@
    Hence, it follows that $y<\infty$ and $k_{2}>\sqrt{k_{1}^{2}+2k_{0}k_{1}+\frac{\lambda k_{1}}{\gamma\alpha}}$ which is again equivalent to $\frac{\lambda k_{1}}{\gamma\alpha}<k_{2}^{2}-k_{1}^{2}-2k_{0}k_{1}$. Thus, we get that $\xi_{3}^{*}>\alpha\hat{\xi}$ holds if and only if $k_{1}=0$ or $\lambda<\gamma\alpha\left(\frac{k_{2}^{2}}{k_{1}}-k_{1}-2k_{0}\right)$ since $k_{2}>0$. Before showing the implication (i), let us also reformulate the equality $\xi_{2}^{*}=\alpha\hat{\xi}$. First, $\xi_{2}^{*}=\alpha\hat{\xi}$ if and only if $\tilde{\xi}_{2}^{*}\geq\alpha\hat{\xi}$ which is equivalent to:
@@
    Since $y<\infty$ (otherwise (A.3) would be wrong), this is equivalent to:
@@
    Now, if $k_{1}=0$ and if $\lambda<\gamma\alpha\left(\frac{k_{2}^{2}}{k_{1}}-k_{1}-2k_{0}\right)$ the inequality is fulfilled since $k_{2}>0$. Thus, statement (i) follows.
    Second, we prove (ii). As in the proof of (i), we start by reformulating the first condition. We note that $\xi_{2}^{*}>\tilde{\alpha}\hat{\xi}$ if and only if $\tilde{\xi}_{2}^{*}>\tilde{\alpha}\hat{\xi}$ which is equivalent to:
@@
    $\lambda(\alpha_{2}k_{2}-\alpha k_{1})>-\gamma\alpha\left(k_{2}^{2}(\tilde{\alpha}-\alpha_{2})+2k_{1}k_{2}\alpha_{2}+2k_{0}k_{2}\alpha_{2}-\alpha k_{1}^{2}-2\alpha k_{0}k_{1}\right).$
    As before, we also reformulate the equality $\xi_{1}^{*}=\tilde{\alpha}\hat{\xi}$. This is indeed true if $\tilde{\xi}_{1}^{*}\geq\tilde{\alpha}\hat{\xi}$. If the term under the square root in the definition of $\tilde{\xi}_{1}^{*}$ (see Theorem 3.2) is 00, then $\tilde{\xi}_{1}^{*}\geq\tilde{\alpha}\hat{\xi}$ is fulfilled since $\frac{2\gamma\tilde{\alpha}^{2}k_{2}}{y}\geq 0$. Now, if this term is positive, $\tilde{\xi}_{1}^{*}\geq\tilde{\alpha}\hat{\xi}$ is equivalent to:
@@
    Now, we get that $\frac{\tilde{\alpha}^{2}-\alpha_{2}^{2}}{\alpha}=\frac{(\tilde{\alpha}-\alpha_{2})(\tilde{\alpha}+\alpha_{2})}{\alpha}=\tilde{\alpha}-\alpha_{2}$ since $\tilde{\alpha}+\alpha_{2}=\alpha$. Hence, the inequality is fulfilled if $\lambda(\alpha_{2}k_{2}-\alpha k_{1})>-\gamma\alpha\left(k_{2}^{2}(\tilde{\alpha}-\alpha_{2})+2k_{1}k_{2}\alpha_{2}+2k_{0}k_{2}\alpha_{2}-\alpha k_{1}^{2}-2\alpha k_{0}k_{1}\right)$ which implies statement (ii) and finishes the proof. ∎
@@
    Using the two properties proven in the proof of Proposition 3.4, we only have to check the boundary values $\xi_{1}^{*}=\tilde{\alpha}\hat{\xi}$ and $\xi_{2}^{*}=\alpha\hat{\xi}$ for continuity. However, this follows immediately by plugging in the values. Now, if $k_{1}>0$, we see that $\hat{X}_{T}>0$ if $\xi^{*}\neq\xi_{3}^{*}$. If $\xi^{*}=\xi_{3}^{*}$, then we note that
@@
$\lim_{y\to 0}f_{1}(y,\lambda)$ $\geq 0\text{ for all $\lambda>C$},$ (A.8)
@@
$\limsup_{\lambda\to\infty}f_{2}(h(\lambda),\lambda)$ $\leq 0.$ (A.11)
@@
Now, if we consider $y^{*}_{\lambda}$ as a function of $\lambda$, it fulfills the assumptions of $h$. Hence, there exists a $\lambda^{*}$ such that $f_{2}(y^{*}_{\lambda^{*}},\lambda^{*})=0$. On the other hand, by the definition of $\lambda\mapsto y_{\lambda}$, we have that $f(y_{\lambda^{*}}^{*},\lambda^{*})=0$, which implies the claim.
@@
    Thus, if $\lambda\leq 2\gamma\alpha(k_{2}-k_{1}-k_{0})$ then it holds that (i) $\Phi(d_{1}(\xi_{3}^{*},0))-\Phi(d_{1}(\alpha\hat{\xi},0))\xrightarrow{y\to 0}1$. On the other hand, if $\lambda>2\gamma\alpha(k_{2}-k_{1}-k_{0})$, then it holds that (ii) $\Phi\left(d_{1}\left(\xi_{1}^{*},0\right)\right)\xrightarrow{y\to 0}1$ as (A.18) entails that $\xi_{1}^{*}\xrightarrow{y\to 0}\infty$. Now, if $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}\leq k_{1}$, it holds that (iii) $\left(k_{0}+k_{1}+\frac{\lambda}{2\gamma\alpha}\right)e^{-\int_{0}^{T}r_{s}\mathrm{d}s}>x_{0}$ for $\lambda>C=-2\gamma\alpha k_{0}$. Furthermore, if $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}\in(k_{1},k_{2}]$, we deduce that (iv) $\left(k_{0}+k_{1}+\frac{\lambda}{2\gamma\alpha}\right)e^{-\int_{0}^{T}r_{s}\mathrm{d}s}>x_{0}$ for $\lambda>C=2\gamma\alpha(x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}-k_{1}-k_{0})$ and, if $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}\leq k_{2}$, it follows directly that (v) $\left(k_{2}+\frac{\lambda}{2\gamma\tilde{\alpha}}-\frac{\alpha}{\tilde{\alpha}}(k_{2}-k_{1}-k_{0})\right)e^{-\int_{0}^{T}r_{s}\mathrm{d}s}>x_{0}$ for $\lambda>2\gamma\alpha(k_{2}-k_{1}-k_{0})(\geq C)$. Finally, if $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}>k_{2}$, then it holds that $C>2\gamma\alpha(k_{2}-k_{1}-k_{0})$ and (vi) $\left(k_{2}+\frac{\lambda}{2\gamma\tilde{\alpha}}-\frac{\alpha}{\tilde{\alpha}}(k_{2}-k_{1}-k_{0})\right)e^{-\int_{0}^{T}r_{s}\mathrm{d}s}>x_{0}$ for $\lambda>C=2\gamma\left(\tilde{\alpha}x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}+\alpha_{2}k_{2}-\alpha(k_{0}+k_{1})\right)$. Note that $\tilde{\alpha}\hat{\xi}\leq\xi_{2}^{*}\leq\alpha\hat{\xi}$ (see (A.16)) and hence (vii) $\xi_{2}^{*}$ shows the same limiting behavior as $\hat{\xi}$ for $y\to 0$. Summarizing these results, we conclude for $\lambda>C$:
@@
      Properties (i), (iii), and (iv) imply:
@@
    Thus, the claim follows. ∎
@@
    Since $y,\gamma,\alpha,\varepsilon>0$ and $k_{1}\geq 0$, we get that $\tilde{\xi}_{3}^{*}>0$ if and only if $(k_{1}+\varepsilon)^{2}>\left(\sqrt{(k_{1}+\varepsilon)^{2}-\varepsilon^{2}}\right)^{2}$, which is equivalent to $\varepsilon^{2}>0$. Thus, the claim, i.e., $\xi_{3}^{*}>0$ for $\lambda>C$.

### v3

@@ 2.2 Optimization Functional
where $X_{s}=x$, $0\leq k_{0},k_{1}\leq k_{2}<\infty$ with $k_{0}+k_{1}\leq k_{2}$, $K_{2}>0$, $0\leq\alpha_{2}<\alpha<\infty$ with $\tilde{\alpha}:=\alpha-\alpha_{2}$. Note that $0<\tilde{\alpha}\leq\alpha$ holds. We also write $F(X_{T})$ instead of $F(0,T,u,x)$ when the trading strategy is clear (for instance, for the optimal terminal wealth $\hat{X}_{T}$), suppressing the initial value.
@@
Since the problem in this paper is a variant of a mean-variance optimization, it inherits also the time-inconsistency of the classical mean-variance approach, see, for instance, Basak and Chabakauri [3] or Björk and Murgoci [5]. There are several ways to approach this time-inconsistency. We take in this paper the perspective of a so-called pre-committed investor. Such an investor optimizes once in the beginning and sticks to this strategy, i.e., the chosen strategy of the investor is not dynamically reoptimized. Note that our problem is not covered by the general theory from Björk and Murgoci [5] which assumes a sophisticated investor, i.e., she/he knows that her/his optimal strategy changes over time and only optimizes over all strategies she/he would actually follow.
@@ 3.1 Derivation of the optimal terminal wealth
    The representation of $\hat{X}_{T}$ also holds in a general complete market (without necessarily using a Black-Scholes market model) if we assume the existence of the multipliers $y$ and $\lambda$. However, that these exist is then not clear. Note that the price density process, $\xi$, in a different complete market would change, but still admit a geometric form. For example, if we replace the Brownian Motion $W$ by a Poisson process $N$ with parameter $\lambda_{2}$ (in the Black-Scholes model), the price density admits a dynamic of the form (see, e.g., Cont and Tankov [11, p.303]):
@@
    The following paragraph considers $L$ as a function of $X$. Obviously, $L$ is not smooth in $k_{1}$ and $k_{2}$, but it is continuous in $[0,\infty)$. Hence, we optimize the piecewise concave function $L$ in the regions $[0,k_{1}]$, $[k_{1},k_{2}]$, and $[k_{2},\infty)$ separately and compare afterwards the maximal points. Now, for the optimization, we get:
@@
    $X_{2}\in[k_{1},k_{2}]$$\Leftrightarrow$ $\xi\in\left[\displaystyle\frac{\lambda\alpha-2\gamma\alpha^{2}(k_{2}-k_{1}-k_{0})}{y},\displaystyle\frac{\lambda\alpha}{y}+\displaystyle\frac{2\gamma\alpha^{2}k_{0}}{y}\right]$
    $\Leftrightarrow$ $\xi\in\left[\alpha\hat{\xi},\bar{\xi}\right],$ (3.6)
@@
    We notice that $\xi_{+}>\bar{\xi}$ and $\max\left\{\alpha\hat{\xi},\xi_{-}\right\}=\xi_{3}^{*}$ where $\xi_{3}^{*}$ is defined as in Theorem 3.2. If $\xi_{-}<\alpha\hat{\xi}$, $X_{2}$ is not in the interval $[k_{1},k_{2})$ which entails that $(\alpha\hat{\xi},\xi_{3}^{*}]=\emptyset$. Otherwise, it holds that $\xi_{-}=\xi_{3}^{*}$ by which we get (3.4) for this case.
@@
    Without loss of generality, we assume that $\hat{\xi}>0$, since otherwise $(0,\tilde{\alpha}\hat{\xi}]=\emptyset$ and there is nothing to show. Then it holds:
@@
    Again, we notice that $\xi_{+}>\tilde{\alpha}\hat{\xi}$ and since the discriminant is non-negative, we get $\xi_{-}=\tilde{\xi}_{1}^{*}$ where we can ignore this case if $\xi_{-}\leq 0$ due to the assumption that $\xi$ has to be in the interval $(0,\tilde{\alpha}\hat{\xi}]$. Now, if the discriminant is negative, then $L(X_{3},y,\xi)\geq-\gamma\alpha^{2}k_{0}^{2}-\lambda\alpha k_{0}$ for all $\xi$ and $\tilde{\xi}_{1}^{*}>\tilde{\alpha}\hat{\xi}$. Hence, $X_{3}$ is optimal for all $\xi\in(0,\tilde{\alpha}\hat{\xi}]$ and it holds that $\xi_{1}^{*}=\tilde{\alpha}\hat{\xi}$ which implies (3.4) for this case.
@@
    Hence, it follows that $X_{4}$ is optimal for $\xi\in(\tilde{\alpha}\hat{\xi},\tilde{\xi}_{2}^{*}]$ which implies (3.4) in this case as well. Hence, (3.4) holds in total.
@@ 4 Numerical Results
In Figure 4, we present the optimal terminal wealth as a function of $\xi_{T}$ for different values of the risk aversion parameter $\gamma$. (Note that the risk aversion parameters are not comparable for the different optimization strategies.) For both the non-protected and the protected case, we compare the optimal terminal wealth stemming from our mean-variance optimization with the EU-optimization for participating life insurance contracts from Lin et al. [24]. To measure the utility, we use, as Lin et al., the following S-shaped utility function: $U(x)=\begin{cases}x^{\tilde{\gamma}}&x\geq 0,\\ -\tilde{\lambda}(-x)^{-\tilde{\gamma}}&x<0\end{cases}$ with $\tilde{\lambda}=2$ and different values for $\tilde{\gamma}$ (Lin et al. [24] used $\tilde{\gamma}=0.5$ in their numerical analysis). Moreover, we added the case that there is no additional surplus participation of the policyholders, i.e., $\alpha_{2}=0$, $k_{0}=0$, and $k_{1}=0$. Both participating optimizations share that the optimal wealth function has three points of significant behavioral change, i.e., at $\xi_{1}^{*}$, $\xi_{2}^{*}$ and $\xi_{3}^{*}$ as in Theorem 3.2. This theorem shows that the optimal wealth decreases before $\xi_{1}^{*}$ and between $\xi_{2}^{*}$ and $\xi_{3}^{*}$, and that the optimal wealth is constant elsewhere. The main difference in the shape of the functions is that we have a piecewise linearity in the mean-variance case but a total non-linearity in the EU case. This effect is already apparent from the respective formulas. Noticeably, the distance between the second and third point of behavioral change is much higher in the EU-optimization than in the mean-variance one. Moreover, the figure shows a moral hazard of the insurance companies offering a non-protected participating life insurance contract since the insurance companies favor a portfolio value of 00 over a portfolio value at the guarantee value level. We can observe this effect for both the mean-variance and the EU-optimization in the figure due to the drop-down of the optimal terminal wealth to zero. Protected products have no such effect when optimizing mean-variance since the insurance company always profits from higher portfolio values. While this effect also occurs when optimizing EU, the probability is smaller since the drop-down is at a far higher value (compared to the non-protected product).
@@
In both, Figure 4 and Figure 5, we can observe a jump of the optimal terminal wealth in the non-protected case, which is typical for problems with option-like constraints, see for instance Basak and Shapiro [4], Chen et al. [8], Nguyen and Stadje [32], Chen et al. [9], or Kraft and Steffensen [20]. The reason is that for certain outcomes, the insurance company’s incentive change and therefore it can gain additional utility by changing to a different regime. [^6] For instance, if the entire wealth for certain outcomes is transferred to the policyholders, the insurance company will target a terminal wealth of zero in these states, in order to gain additional wealth in states where it benefits more. (Note that the overall expectation is fixed due to the budget constraint.) Moreover, the piecewise linearity observed also occurs in, e.g., Avanzi et al. [1] who considered a minimization of a quadratic function under solvency constraints including option-like constraints as the expected shortfall. This piecewise linearity is typical for problems with some kind of quadratic penalties, whereas the option-like constraints entail the existence of the plateaus and the jumps.
@@ Appendix A Proofs
    We start with the proof of (i). First, if $\hat{\xi}=0$, then it holds that $\xi_{2}^{*}=0$ by definition, i.e., we get the claim. Hence, we assume that $\hat{\xi}>0$. Next, we reformulate the condition $\xi_{3}^{*}>\alpha\hat{\xi}$ to get a condition for $\lambda$. When plugging in the formulas as defined in Theorem 3.2, it holds that $\xi_{3}^{*}>\alpha\hat{\xi}$ if and only if
@@
    Hence, it follows that $y<\infty$ and $k_{2}>\sqrt{k_{1}^{2}+2k_{0}k_{1}+\frac{\lambda k_{1}}{\gamma\alpha}}$ which is again equivalent to $\frac{\lambda k_{1}}{\gamma\alpha}<k_{2}^{2}-k_{1}^{2}-2k_{0}k_{1}$. Thus, we get that $\xi_{3}^{*}>\alpha\hat{\xi}$ holds if and only if $k_{1}=0$ or $\lambda<\gamma\alpha\left(\frac{k_{2}^{2}}{k_{1}}-k_{1}-2k_{0}\right)$ since $k_{2}>0$. Before showing the implication (i), let us also reformulate the equality $\xi_{2}^{*}=\alpha\hat{\xi}$, which we want to show. First, $\xi_{2}^{*}=\alpha\hat{\xi}$ if and only if $\tilde{\xi}_{2}^{*}\geq\alpha\hat{\xi}$ which is equivalent to:
@@
    Since $y<\infty$ (otherwise (A.3) would not hold), this is equivalent to:
@@
    Now, if $k_{1}=0$ or if $\lambda<\gamma\alpha\left(\frac{k_{2}^{2}}{k_{1}}-k_{1}-2k_{0}\right)$ the inequality is fulfilled since $k_{2}>0$. Thus, statement (i) follows.
    Second, we prove (ii). By defintion, it holds that $\hat{\xi}>0$ in order to get $\xi_{2}^{*}>\tilde{\alpha}\hat{\xi}$. As in the proof of (i), we start by reformulating the first condition. We note that $\xi_{2}^{*}>\tilde{\alpha}\hat{\xi}$ if and only if $\tilde{\xi}_{2}^{*}>\tilde{\alpha}\hat{\xi}$ which is equivalent to:
@@
    $\lambda(\alpha_{2}k_{2}-\alpha k_{1})>-\gamma\alpha\left(k_{2}^{2}(\tilde{\alpha}-\alpha_{2})+2k_{1}k_{2}\alpha_{2}+2k_{0}k_{2}\alpha_{2}-\alpha k_{1}^{2}-2\alpha k_{0}k_{1}\right).$ (A.4)
    As before, we also reformulate the equality $\xi_{1}^{*}=\tilde{\alpha}\hat{\xi}$. This equality is indeed true if $\tilde{\xi}_{1}^{*}\geq\tilde{\alpha}\hat{\xi}$. If the term under the square root in the definition of $\tilde{\xi}_{1}^{*}$ (see Theorem 3.2) is 00, then $\tilde{\xi}_{1}^{*}\geq\tilde{\alpha}\hat{\xi}$ is fulfilled since $\frac{2\gamma\tilde{\alpha}^{2}k_{2}}{y}\geq 0$. Now, if this term is positive, $\tilde{\xi}_{1}^{*}\geq\tilde{\alpha}\hat{\xi}$ is equivalent to:
@@
    Now, we get that $\frac{\tilde{\alpha}^{2}-\alpha_{2}^{2}}{\alpha}=\frac{(\tilde{\alpha}-\alpha_{2})(\tilde{\alpha}+\alpha_{2})}{\alpha}=\tilde{\alpha}-\alpha_{2}$ since $\tilde{\alpha}+\alpha_{2}=\alpha$. Hence, the inequality is fulfilled if $\lambda(\alpha_{2}k_{2}-\alpha k_{1})>-\gamma\alpha\left(k_{2}^{2}(\tilde{\alpha}-\alpha_{2})+2k_{1}k_{2}\alpha_{2}+2k_{0}k_{2}\alpha_{2}-\alpha k_{1}^{2}-2\alpha k_{0}k_{1}\right)$ which by equation (A.4) implies statement (ii) and finishes the proof. ∎
@@
    Using the two properties proven in the proof of Proposition 3.4, we only have to check the boundary values $\xi_{1}^{*}=\tilde{\alpha}\hat{\xi}$ and $\xi_{2}^{*}=\alpha\hat{\xi}$ for continuity. However, this follows immediately by plugging in the values. Now, if $k_{1}>0$, we see that $\hat{X}_{T}>0$ if $\xi^{*}\neq\xi_{3}^{*}$. If $\xi^{*}=\xi_{3}^{*}$, by the definition of $\xi^{*}$ in Theorem 3.2, it actually holds that $\xi_{3}^{*}>\alpha\hat{\xi}$. We note that
@@
$\lim_{y\to 0}f_{1}(y,\lambda)$ $>0\text{ for all $\lambda>C$},$ (A.9)
@@
$\lim_{\lambda\to\infty}f_{2}(h(\lambda),\lambda)$ ==$\infty$.= - ∞ . (A.12)
@@
Now, if we consider $y^{*}_{\lambda}$ as a function of $\lambda$, it fulfills the assumptions of $h$. Hence, there exists a $\lambda^{*}$ such that $f_{2}(y^{*}_{\lambda^{*}},\lambda^{*})=0$. On the other hand, by the definition of $\lambda\mapsto y_{\lambda}$, we have that $f_{1}(y_{\lambda^{*}}^{*},\lambda^{*})=0$, which implies the claim.
@@
    Thus, if $\lambda\leq 2\gamma\alpha(k_{2}-k_{1}-k_{0})$ then it holds that (i) $\Phi(d_{1}(\xi_{3}^{*},0))-\Phi(d_{1}(\alpha\hat{\xi},0))\xrightarrow{y\to 0}1$ and $\Phi(d_{1}(\xi_{1}^{*},0))\xrightarrow{y\to 0}0$. On the other hand, if $\lambda>2\gamma\alpha(k_{2}-k_{1}-k_{0})$, then it holds that (ii) $\Phi\left(d_{1}\left(\xi_{1}^{*},0\right)\right)\xrightarrow{y\to 0}1$ as (A.19) entails that $\xi_{1}^{*}\xrightarrow{y\to 0}\infty$. Now, if $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}\leq k_{1}$, it holds that (iii) $\left(k_{0}+k_{1}+\frac{\lambda}{2\gamma\alpha}\right)e^{-\int_{0}^{T}r_{s}\mathrm{d}s}>x_{0}$ for $\lambda>C=-2\gamma\alpha k_{0}$. Furthermore, if $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}\in(k_{1},k_{2}]$, we deduce that (iv) $\left(k_{0}+k_{1}+\frac{\lambda}{2\gamma\alpha}\right)e^{-\int_{0}^{T}r_{s}\mathrm{d}s}>x_{0}$ for $\lambda>C=2\gamma\alpha(x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}-k_{1}-k_{0})$ and, if $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}\leq k_{2}$, it follows directly that (v) $\left(k_{2}+\frac{\lambda}{2\gamma\tilde{\alpha}}-\frac{\alpha}{\tilde{\alpha}}(k_{2}-k_{1}-k_{0})\right)e^{-\int_{0}^{T}r_{s}\mathrm{d}s}>x_{0}$ for $\lambda>2\gamma\alpha(k_{2}-k_{1}-k_{0})(\geq C)$. Finally, if $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}>k_{2}$, then it holds that $C>2\gamma\alpha(k_{2}-k_{1}-k_{0})$ and (vi) $\left(k_{2}+\frac{\lambda}{2\gamma\tilde{\alpha}}-\frac{\alpha}{\tilde{\alpha}}(k_{2}-k_{1}-k_{0})\right)e^{-\int_{0}^{T}r_{s}\mathrm{d}s}>x_{0}$ for $\lambda>C=2\gamma\left(\tilde{\alpha}x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}+\alpha_{2}k_{2}-\alpha(k_{0}+k_{1})\right)$. Note that $\tilde{\alpha}\hat{\xi}\leq\xi_{2}^{*}\leq\alpha\hat{\xi}$ (see (A.17)) and hence (vii) $\xi_{2}^{*}$ shows the same limiting behavior as $\hat{\xi}$ for $y\to 0$. Summarizing these results, we conclude for $\lambda>C$:
@@
      Properties (i), (iii), (iv), $\lambda>-2\gamma\alpha k_{0}$, and (A.17) imply:
@@
    Note that since either $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}>k_{2}$, or $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}\leq k_{2}$ and $\lambda>2\gamma\alpha(k_{2}-k_{1}-k_{0})$ or $\lambda\in(C,2\gamma\alpha(k_{2}-k_{1}-k_{0})]$, we always have in (a) or in (d) ”$>x_{0}$”. Thus, the claim follows. ∎
@@
    Since $y,\gamma,\alpha,\varepsilon>0$ and $k_{1}\geq 0$, we get that $\tilde{\xi}_{3}^{*}>0$ if and only if $(k_{1}+\varepsilon)^{2}>\left(\sqrt{(k_{1}+\varepsilon)^{2}-\varepsilon^{2}}\right)^{2}$, which is equivalent to $\varepsilon^{2}>0$. Thus, the claim, i.e., $\xi_{3}^{*}>0$ for $\lambda>C$, follows.